Perimeter of a rectangle is 180 ft. the width is 1/2 of the length what is the length?

Answer by checkley77(12844)

W=L/2
P=2L+2W
180=2L+2*L/2
180=2L+L
3L=180
L=180/3
L=60 ANS.
W=60/2=30 ANS.
PROOF:
180=2*60+2*30
180=120+60
180=180